A baseball dugout has a height of 8 ft and the volume of 1080 cubic feet
And the answer choice are
17ft by 7 ft
16ft by 8 ft
15ft by 9 ft
14 ft by 10 ft and how to get the answers

Answer by math_helper(2461)


The Volume V, is L*W*H: V = L*W*H
Re-arrange the above to L*W = V/H

V is given as 1080 cu ft and H as 8 ft. V/H = L*W = 135 ==> you must find the pair of numbers that multiply to 135 sq ft.